OpenStudy (mgriffin16):

One side of a square is 5n inches long. Another side is n + 16 inches long. How long is a side of the square?

OpenStudy (welshfella):

all sides of a square are equal therefore 5n = n + 16 Solve for n

OpenStudy (wolframwizard):

Since the sides of a square are all equal in length we know 5n=n+16. Subtract n from both sides to get 4n=16 and then divide by 4 to get n=4. This is not the length of the side though. Plug 4 back in to either expression to get each side is 20 units.5*4=4+16=20

OpenStudy (mgriffin16):

so would it be 16 or 20?

OpenStudy (wolframwizard):

20